OpenStudy (anonymous):

Solve each equation check your answer r-17=7

OpenStudy (anonymous):

r-17=7 (original equation) (r-17) +17= (7 )+ 17 (add 17 to to both sides to get r by itself) r + 0 = 24 r = 24

OpenStudy (anonymous):

thank you for help :)